College Physics

An ideal monooatomic gas expands isothermally from 0.500m^3 to1.25m^3 at a constant temp. of 675 K.If the intial pressureis 1.00x10^5 Pa, what is (a) thework done on the gas, (b) thethermal energy transfer Q, and (c) the change in the internalenergy.

Answers:
a) -4.58x10^4 Jb) 4.58x10^4 Jc) 0
Answer #1
The ideal monoatomic gas expands isothermally from 0.500m3 to 1.25 m3 at a constant temperature of T= 675 KThe initial pressure is P1= 1.00 * 105Pa(a)The work done on the gas isW = P1 * ΔVor W = 1.00 * 105 * (1.25 - 0.500) = 0.75 *105 J = 7.5 * 104 J(b)The thermal energy transfer Q isQ = n * R * THere,n = 1,R = 8.314 J/mol/Kor Q = 1 * 8.314 * 675 = 5611.95 J = 0.56 * 104J(c)The change in the internal energy isU = W - Q = (7.5 * 104 - 0.56 * 104) =6.94 * 104 J